OpenStudy (anonymous):

Find the product of (x + 5)(x - 5). (1 point) x2 - 10x + 25 x2 + 10x + 25 x2 - 25 x2 + 25 Please help asap!!!!!

Vocaloid (vocaloid):

hint: (x+A)(x-A) = x^2 - A^2 let A = 5

OpenStudy (anonymous):

okay..

Vocaloid (vocaloid):

(x+A)(x-A) = x^2 - A^2 (x+5)(x-5) = ?

OpenStudy (anonymous):

Another way to think about this:\[(x+5)(x-5)=x(x-5)+5(x-5)\]by the distributive property. Do you think you can simplify it now?

OpenStudy (lalaland_lauren):

The answer is C.

OpenStudy (anonymous):

okay that was my final solution

OpenStudy (anonymous):

thanks!!